# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 4.1.2 Fuel load
Another crucial variable in fuel conditions is fuel load, which refers to the total biomass per unit area and provides a baseline for the available biomass in wildfires (McNorton and Di Giuseppe, 2024). Fuel load influences the severity and spread of fires (García-Llamas et al., 2019; Alonso-Rego et al., 2020; McNorton and Di Giuseppe, 2024). It can be categorized into live fuel load and dead fuel load, or further divided by size into fine fuel load—such as grass, small shrubs, twigs, and leaves—and coarse fuel load, including trunks and logs (Gould et al., 2011). Moreover, the definition of fuel load varies slightly across different studies. For instance, Lee et al. (2022) defines fuel load as the amount of fuel in the upper parts of trees, specifically canopy fuel load, while McNorton and Di Giuseppe (2024) defines the above-ground component of fuel load as above-ground biomass. Notably, there is an emphasis on dead fine fuel because it is the primary carrier of fire ignition and spread (Cruz et al., 2003; Elia et al., 2020; D’Este et al., 2021), while canopy fuel load is highlighted because crown fires typically consume fine fuel, resulting in more intense and faster-spreading fires (Arellano-Pérez et al., 2018; Walker et al., 2020; de Groot et al., 2022).
In addition to fuel load estimation using physical or semi-empirical models based on remote sensing data, empirical model-based methods are simpler and more commonly used. For instance, Arellano-Pérez et al. (2018) used the random forest method and Sentinel-2 multispectral and vegetation index data to estimate the canopy fuel load of Pinus radiata and Pinus pinaster stands in northwestern Spain. However, due to the low penetration capability of optical sensors, the accuracy of canopy fuel load estimation in this study was limited. Furthermore, Li et al. (2021a) compared the effectiveness of L-band SAR (ALOS-PALSAR) and Landsat 7 ETM+ in estimating fuel load in coniferous mixed forests and found that L-band SAR outperformed optical data in all three fuel load types, with estimation accuracy further improved when both were integrated. Li et al. (2022b) used a data-driven random forest model to correlate long-term spatiotemporal characteristics derived from Landsat 8 OLI and Sentinel-1 data with field-measured fuel load, estimating the fine fuel load (FFL) of pine forests in some areas of Sichuan Province, China. The study found that combining SAR and multispectral data improved FFL estimation accuracy, and temporal features were more important than spatial features.
In addition, the more recent NASA Global Ecosystem Dynamics Investigation (GEDI) full-waveform LiDAR sensor has been widely used for large-scale above-ground biomass estimation and vegetation canopy structure analysis (Myroniuk et al., 2023; Liang et al., 2023; Aragoneses et al., 2024; Guerra-Hernández et al., 2024), aiding wildfire simulation, assessment of wildfire impacts on vegetation, and understanding wildfire severity. However, only a few studies have directly used GEDI LiDAR data to estimate fuel load. For example, Leite et al. (2022) used unmanned aerial vehicle (UAV) simulations and GEDI data to study the ability to stratify fuel load in grasslands, savannas, and forests within tropical savannas in Brazil. The study found that GEDI performed poorly in estimating fuel load for ground and low-lying vegetation, and was significantly affected by terrain.
@@
Finally, it is worth noting that due to the similarity in definitions, particularly the emphasis on crown fires, some studies have used above-ground biomass as a proxy for fuel load (Eames et al., 2021; Vega et al., 2022). Additionally, other studies have used different proxies for fuel load. For instance, Su et al. (2021) used fractional vegetation cover (FVC) as a proxy for fuel load in a study on wildfire drivers in tropical rainforests in southern China. FVC represents the percentage of an area covered by green vegetation within a statistical region. In practical applications, the MODIS NDVI product is commonly used as a proxy for FVC. Helman et al. (2015) analyzed time-series NDVI data from MODIS to distinguish between woody and herbaceous vegetation in Mediterranean forest regions, using long-term mean woody NDVI and its trends as a proxy for fire fuel. In actual wildfire risk prediction, using proxies instead of directly estimating fuel load is more common because it is simpler and avoids cumulative errors.
@@ 4.1.3 Fuel types
Another method to describe fuel properties, aside from fuel moisture code and fuel load, is by categorizing fuels into homogeneous groups that exhibit similar characteristics under specified burning conditions. These groups are defined as "fuel types" or "fuel models" (Arroyo et al., 2008). The description of fuel types is generally complex, taking into account various fuel characteristics such as vegetation type, canopy height, canopy base height, canopy bulk density, vegetation cover ratio, horizontal and vertical continuity, fuel load, fuel moisture, and biomass (Chuvieco et al., 2003; Pettinari and Chuvieco, 2020; Aragoneses and Chuvieco, 2021). In recent years, various standardized fuel classification systems have been proposed in different regions (Table 1), including the NFDRS (Cohen, 1985), Northern Forest Fire Laboratory (NFFL) (Albini, 1976; Burgan, 1984), Fuel Characteristics Classification System (FCCS) (Sandberg et al., 2001; Ottmar et al., 2007), McArthur fuel types (McArthur, 1966, 1967), Canadian Fire Behaviour Prediction (FBP) (Group et al., 1992), and Prometheus (Riaño et al., 2002).
@@
Similar to the assessment methods for fuel load and fuel moisture, fuel type mapping can also be conducted through field sampling and remote sensing techniques, including the use of optical, microwave, and LiDAR sensors (García et al., 2012; Pettinari and Chuvieco, 2020; Aragoneses and Chuvieco, 2021). This generally involves the classification of vegetation types and the retrieval of vegetation structure and biochemical parameters such as canopy, followed by the classification of fuel types using expert knowledge or experimental data (Chuvieco et al., 2020). At large scales, such as global or continental levels, fuel type maps are usually created by combining satellite imagery with land use types and the aforementioned fuel classification systems (Pettinari et al., 2013; Pettinari and Chuvieco, 2016; Aragoneses and Chuvieco, 2021).
In addition to detailed classification of fuel types, many wildfire risk assessment studies based on multi-source remote sensing data employ simpler, but coarser, land use type maps as proxies for fuel types to analyze the impact of different land cover types or human activities on wildfire risk. For instance, Jaafari et al. (2019) used logistic regression to analyze the impact of various explanatory variables, including topography, elevation, TWI, meteorological factors, and land use, on wildfire risk in the Zagros ecoregion of western Iran. They found that the most important variable was land use type, with dry farmland and forests being the most susceptible to wildfires. Additionally, the type and continuity of vegetation determined the likelihood of large fires. Similarly, Ager et al. (2014b) analyzed wildfire occurrence and size on the islands of Sardinia, Italy, and Corsica, France, finding that early in the fire season, the probability of large fires was ten times higher in agricultural lands in southern Sardinia compared to forest and shrubland land use types near roads. However, by the end of the fire season, the latter land use types exhibited a higher probability of fires. Conversely, Calviño-Cancela et al. (2016) analyzed the wildfire risk in the wildland-urban interface (WUI) and LULC in northern Spain, finding that primary forests and agricultural areas had the lowest fire risk. Similarly, Donovan et al. (2020) assessed wildfire risk in the Great Plains ecoregion of the United States, considering five land use types: grasslands, woody vegetation, croplands, pastures and hayfields, and developed areas. Their study found that woody vegetation and grasslands were most prone to wildfires and large fires, while other land use types had lower risks. Among them, croplands had the lowest probability of large fires, and as the percentage of crop cover increased, the likelihood of large fires decreased.
@@ 4.1.4 Fuel continuity
In addition to vertical continuity, horizontal fuel continuity is also crucial. However, there are various definitions of horizontal continuity. For example, Frandsen et al. (1979) defines horizontal continuity as the degree of variation in the physical characteristics of fuels within a given area, with an emphasis on the impact of fuel type. Conversely, a more commonly used definition in describing horizontal continuity focuses on the spacing between fuels (Ritter et al., 2023). Tree spacing significantly affects the intensity and scale of wildfires (Kim et al., 2016). For instance, Qadir et al. (2021) showed in a study on wildfire risk prediction in Nepal that areas with high forest cover density have a higher risk of fires. In contrast, when there is canopy discontinuity, fire spread may slow down or even be extinguished (Taneja et al., 2021). For example, Kim et al. (2016) used the Fire Dynamics Simulator (WFDS) to model fire behavior in Korean pine (Pinus densiflora) and found that when the spacing between trees is 6 meters or more, most crown fires stop spreading within 100 meters. Therefore, fuel management techniques that fragment fuels can be used to suppress wildfires (Srivastava et al., 2014; Harrison et al., 2021). Based on our review, research on horizontal fuel continuity is less prevalent compared to vertical continuity, likely because many studies use proxies rather than directly assessing horizontal continuity. For instance, many crown fire spread models use canopy bulk density to describe horizontal fuel continuity (Wagner, 1993; Cruz et al., 2005). Additionally, proxies such as fraction vegetation cover, canopy cover, vegetation cover density, and various vegetation indices are also commonly used.
In summary of fuel conditions, it is noteworthy that various fuel state description parameters are generally used in studies evaluating their effectiveness in wildfire risk modeling but are rarely directly utilized in wildfire risk assessment models. Commonly used wildfire risk assessment models often directly employ raw multispectral or microwave data or indices derived from these data. This could be due to several reasons:
@@
Firstly, evaluating fuel state is complex, often requiring multiple sensors, and the models built for this purpose tend to have poor generalization ability, making them suitable only for specific regions. Additionally, the availability of sensors is a limiting factor for the application of certain fuel state description parameters. For example, estimating vertical canopy structure typically requires LiDAR data. However, satellite-based LiDAR canopy structure is limited by spatiotemporal resolution, and airborne LiDAR has limited coverage and high costs.
@@
Moreover, most wildfire risk assessment studies require long time series data, which, due to cost constraints, are often based on low-cost and readily available data such as multispectral or microwave satellite data. These data, due to limitations in spatiotemporal resolution and canopy penetration, often struggle to accurately retrieve various descriptive parameters. The descriptive parameters obtained from inaccurate data inversions may lead to error accumulation, which can further impact prediction accuracy.
Lastly, data-driven wildfire risk prediction algorithms are becoming increasingly popular, and these algorithms can either rely on or bypass manual feature engineering, allowing them to directly learn the necessary abstractions from features or raw data.
@@ 4.2 Weather and climate conditions
Weather and climate factors are major drivers of wildfire activity, influencing the occurrence, burned area, and fire behavior (Swetnam and Betancourt, 1990; Bessie and Johnson, 1995; Abatzoglou and Kolden, 2013). Weather conditions are central to almost all global fire danger risk systems. High temperatures, low relative humidity, insufficient precipitation, and strong winds have been proven to be key factors in the ignition of fires in the short term (Keane et al., 2001; Zacharakis and Tsihrintzis, 2023b), and they also affect fire size and spread (Abatzoglou and Kolden, 2011). On the other hand, climate conditions generally refer to the long-term weather conditions of a region and their variations. Some studies specifically define these as the collective atmospheric conditions months or even years before the start of the fire season, emphasizing the use of prior climate conditions to convert fuel conditions into wildfire potential (Abatzoglou and Kolden, 2013). This approach enables predictions of large-scale fire risks over seasonal, annual, or even multi-decadal time scales.
@@ 4.2.2 Climate conditions
Climate conditions primarily refer to annual or multi-year meteorological parameters such as temperature, humidity, precipitation, and evaporation. Unlike short-term meteorological conditions that affect fuel status temporarily, climate conditions play a crucial role in determining fuel types, distribution, life cycles, and other factors that influence fuel characteristics, such as fuel load and continuity (Aldersley et al., 2011; Jaafari et al., 2017; Tavakkoli Piralilou et al., 2022). These fuel characteristics respond to climatic patterns accumulated before the start of the fire season (Chen et al., 2016). Consequently, climate variables are often the primary drivers of long-term wildfire activity (Pollina et al., 2013; Jain et al., 2020). Furthermore, climate conditions also explain strong seasonal variations in wildfire risk in certain regions (Martell et al., 1989), particularly in the context of intensified climate change in recent years, which has raised concerns about its long-term impact on the occurrence of large-scale wildfires (Prior and Eriksen, 2013; Jolly et al., 2015; Jain et al., 2024).
@@
Climate conditions are widely used in wildfire risk prediction. Specifically, Chen et al. (2016) used large-scale sea surface temperature anomalies to predict the severity of the fire season in South America on a seasonal time scale. Their study found that using a single ocean climate index could predict about 48% of the burned area globally for up to three months or longer before the peak burning month. In a study by Jaafari et al. (2019) in the Zagros ecoregion of western Iran, they used three climate conditions: annual temperature, rainfall, and wind effects. However, the results showed that only the annual temperature parameter was relatively important. Similarly, Naderpour et al. (2020) studied the northern beaches of New South Wales, Australia, using three climate factors: annual rainfall, annual humidity, and annual wind speed. Holdrege et al. (2024) studied the Great Basin sagebrush regions across 13 western U.S. states and used three climate factors: annual mean temperature, annual precipitation, and the proportion of summer (June to August) precipitation to total annual precipitation. The study found that areas with low summer precipitation proportion, medium to high annual precipitation, and high temperatures had the highest wildfire probability. Similarly, Aldersley et al. (2011) used annual cumulative precipitation, annual lightning frequency, monthly precipitation frequency, monthly average temperature, and climate data from the British Atmospheric Data Centre (BADC) to study the drivers of monthly fire area on a global and regional scale. They found that climate factors such as high temperatures, moderate precipitation, and drought were the most important drivers globally. Mansuy et al. (2019) conducted a study across the U.S. and Canada and found that climate variables were the primary control factors for the burned area in most ecoregions, both inside and outside protected areas, surpassing landscape and anthropogenic factors. The climate data used in their study included annual average precipitation, annual average relative humidity, the Hargreaves climate moisture deficit (an annual measure of energy and moisture), and degree days below 18°C. Abdollahi and Pradhan (2023) conducted a wildfire prediction study in the Gippsland region of Victoria, Australia, and found that humidity, wind speed, and rainfall were the most influential factors in wildfire prediction. Zhao and Liu (2021) used data from over 600 meteorological stations in forest, grassland, and shrubland areas to calculate two drought indices, the Keetch-Byram Drought Index and the Standardized Precipitation Index. They then used global-scale PDSI data to mask these indices and assess long-term wildfire risk in China. The results showed that drought indices had a stronger correlation with seasonal and annual fire occurrences compared to climate indicators such as temperature and precipitation.

### v5

@@ 4.1.2 Fuel Load
Fuel load refers to the amount of combustible biomass available per unit area which influences wildfire intensity and spread (McNorton and Di Giuseppe, 2024). It can be classified as live or dead, or by size into fine fuels (e.g., grass, shrubs, twigs) and coarse fuels (e.g., trunks, logs) (Gould et al., 2011). Dead fine fuel plays a critical role in fire ignition and spread, while canopy fuel load fuels crown fires, increasing both fire intensity and spread rate (Cruz et al., 2003; Arellano-Pérez et al., 2018; Walker et al., 2020). Fuel load measurements are typically obtained through direct sampling, physical models, or remote sensing techniques. The integration of SAR, LiDAR, and multispectral data has significantly enhanced the accuracy of fuel load estimation.
In practical wildfire risk prediction, directly estimating fuel load is often challenging due to its inherent complexity and data limitations, which may lead to inaccurate estimations. Consequently, the use of proxies rather than direct fuel load estimation is more common in wildfire risk prediction, as it simplifies the process and avoids the accumulation of errors.
@@
First, due to its definitional similarity to fuel load, particularly its emphasis on above-ground biomass, some studies have employed above-ground biomass as a proxy for fuel load (Eames et al., 2021; Vega et al., 2022; McNorton and Di Giuseppe, 2024). Additionally, other studies have utilized alternative proxies for fuel load. For instance, Su et al. (2021), in their study on wildfire driving factors in certain tropical rainforest regions of southern China, used fractional vegetation cover (FVC) as a proxy for fuel load. FVC represents the percentage of an area covered by green vegetation relative to the total statistical area. In practical applications, MODIS NDVI products are often used as proxies for FVC. Furthermore, Helman et al. (2015) applied time-series analysis to MODIS NDVI data to distinguish between woody and herbaceous vegetation in Mediterranean forest regions. Their findings suggested that the long-term mean woody NDVI and its trend could serve as alternative indicators for wildfire fuel assessment.
@@ 4.1.3 Fuel Types
Fuel types categorize fuels based on shared characteristics, such as vegetation type, canopy height, fuel load, and biomass, which influence wildfire behavior under specific conditions (Arroyo et al., 2008). Classification systems for fuel types vary globally. For instance, the U.S. National Fire Danger Rating System (NFDRS), Northern Forest Fire Laboratory (NFFL), and Canada’s Fire Behaviour Prediction (FBP) system are widely used for wildfire risk prediction and management (Cohen, 1985; Albini, 1976; Group et al., 1992).
Table 1 presents a comparison of major fuel classification systems. The NFDRS includes 20 fuel types, NFFL defines 13 types, and the U.S. Fuel Characteristics Classification System (FCCS) has 216 fuel types. In Australia, McArthur’s system categorizes fuel into three types, while the FBP system in Canada uses 16 fuel types. The Prometheus system, applied in Mediterranean Europe, defines seven types (Arroyo et al., 2008).
@@
In addition to the detailed classification of fuel types, most wildfire risk assessment studies based on multi-source remote sensing data utilize simplified LULC maps as proxies for fuel type to analyze the influence of different land cover types and human activities on wildfire risk. For instance, Jaafari et al. (2019) employed logistic regression to assess the influence of various explanatory variables—including topography, elevation, Topographic Wetness Index (TWI), meteorological factors, and land use—on wildfire risk in the Zagros ecological zone of western Iran. The study identified land use as the most significant factor, with dry farmland and forests being the areas most susceptible to wildfires. Moreover, the type and continuity of vegetation were found to be critical determinants of large fire occurrences.
Similarly, Ager et al. (2014b) analyzed wildfire occurrence and scale in Sardinia (Italy) and Corsica (France). Their research revealed that, at the beginning of the the fire season, the probability of large fires occurring in agricultural lands in southern Sardinia was ten times higher than in forested and shrubland areas near roads. However, as the fire season progressed, the latter exhibited a higher wildfire probability. Conversely, Calviño-Cancela et al. (2016) investigated the relationship between WUI and LULC in northern Spain and found that primary forests and agricultural areas exhibited the lowest wildfire risk.
Likewise, Donovan et al. (2020) considered five land use types—grasslands, woody vegetation, croplands, rangelands and hayfields, and developed areas—in an assessment of wildfire occurrence risk in the Great Plains ecoregion of the United States. Their findings indicated that woody vegetation and grasslands were most prone to both wildfire ignition and large fires, whereas other land use types exhibited lower susceptibility. Among them, croplands had the lowest probability of large fires, and as crop coverage increased, the likelihood of large fire occurrences decreased.
@@ 4.1.4 Fuel Continuity
In addition to vertical continuity, horizontal fuel continuity is also crucial. There are various definitions of horizontal continuity. For example, Frandsen et al. (1979) defines horizontal continuity as the degree of variation in the physical characteristics of fuels within a given area, with an emphasis on the impact of fuel type. Conversely, a more commonly used definition in describing horizontal continuity focuses on the spacing between fuels (Ritter et al., 2023). Tree spacing significantly affects the intensity and scale of wildfires (Kim et al., 2016). For instance, Qadir et al. (2021) showed in a study on wildfire risk prediction in Nepal that areas with high forest cover density have a higher risk of fires. In contrast, when there is canopy discontinuity, fire spread may slow down or even be extinguished (Taneja et al., 2021). For example, Kim et al. (2016) used the Fire Dynamics Simulator (WFDS) to model fire behavior in Korean pine (Pinus densiflora) and found that when the spacing between trees is 6 meters or more, most crown fires stop spreading within 100 meters. Therefore, fuel management techniques that fragment fuels can be used to suppress wildfires (Srivastava et al., 2014; Harrison et al., 2021).
Despite the demonstrated importance of horizontal fuel continuity, our review indicates that it has received comparatively less attention than vertical continuity. This discrepancy is likely due to the frequent use of proxies rather than direct measurements. For example, many crown fire spread models rely on canopy bulk density to approximate horizontal fuel continuity (Wagner, 1993; Cruz et al., 2005). Additional proxies, such as fractional vegetation cover, canopy cover, vegetation cover density, and various vegetation indices, are also commonly employed.
In summary, fuel conditions play a crucial role in wildfire risk modeling. However, directly using fuel condition descriptors derived from remote sensing or multi-source geospatial data in large-scale and long-term wildfire risk prediction models is uncommon. Instead, as illustrated in Figure 3, most models rely on raw multispectral or microwave data and meteorological parameters, or indices derived from these sources, which have been shown to correlate with fuel conditions. This practice is driven by several considerations:
@@
First, accurately assessing fuel conditions is inherently complex and often requires integrating data from multiple sensors. Models designed for this purpose frequently exhibit poor generalization, making them applicable only to specific regions or temporal scales. Additionally, the availability of certain sensors poses a limitation. For instance, estimating vertical canopy structure relies heavily on LiDAR data. However, satellite-based LiDAR canopy measurements are constrained by spatial and temporal resolution, while airborne LiDAR is limited by coverage and high costs.
@@
Second, most wildfire risk assessment studies necessitate long-term time-series data, which are typically sourced from low-cost, widely available datasets such as multispectral or microwave satellite data. Due to constraints in spatial resolution and canopy penetration, these data often fail to accurately retrieve detailed fuel condition descriptors. Consequently, inaccuracies in inversion processes may lead to cumulative errors, further impacting prediction accuracy.
Finally, the growing popularity of data-driven wildfire risk prediction algorithms reduces the reliance on explicitly defined fuel condition descriptors. These algorithms, capable of bypassing manual feature engineering, directly learn the necessary abstractions from raw data or features. This adaptability minimizes the need for explicit fuel condition parameters while maintaining high predictive performance.
@@ 4.2 Weather and Climate Conditions
Weather and climate factors are major drivers of wildfire activity, influencing the occurrence, burned area, and fire behavior (Swetnam and Betancourt, 1990; Bessie and Johnson, 1995; Abatzoglou and Kolden, 2013). Weather conditions are central to almost all global fire danger risk systems. High temperatures, low relative humidity, insufficient precipitation, and strong winds have been proven to be key factors in the ignition of fires in the short term (Keane et al., 2001; Zacharakis and Tsihrintzis, 2023), and they also affect fire size and spread (Abatzoglou and Kolden, 2011). On the other hand, climate conditions generally refer to the long-term weather conditions of a region and their variations. Some studies specifically define these as the collective atmospheric conditions months or even years before the start of the fire season, emphasizing the use of prior climate conditions to convert fuel conditions into wildfire potential (Abatzoglou and Kolden, 2013). This approach enables predictions of large-scale fire risks over seasonal, annual, or even multi-decadal time scales.
@@ 4.2.2 Climate Conditions
Climate conditions primarily refer to annual or multi-year meteorological parameters such as temperature, humidity, precipitation, and evaporation. Unlike short-term meteorological conditions that affect fuel status temporarily, climate conditions play a crucial role in determining fuel types, distribution, life cycles, and other factors that influence fuel characteristics, such as fuel load, type, and continuity (Aldersley et al., 2011; Jaafari et al., 2017; Tavakkoli Piralilou et al., 2022). These fuel characteristics respond to climatic patterns accumulated before the start of the fire season (Chen et al., 2016). Consequently, climate variables are often the primary drivers of long-term wildfire activity (Pollina et al., 2013; Jain et al., 2020). Furthermore, climate conditions explain strong seasonal variations in wildfire risk in certain regions (Martell et al., 1989). Intensified climate change in recent years has raised concerns about its long-term impacts on large-scale wildfires (Prior and Eriksen, 2013; Jolly et al., 2015; Jain et al., 2024).
@@
Climate conditions are widely used in wildfire risk prediction. Specifically, Chen et al. (2016) used large-scale sea surface temperature anomalies to predict the severity of the fire season in South America on a seasonal time scale. Their study found that using a single ocean climate index could predict about 48% of the burned area globally for up to three months or longer before the peak burning month. In a study by Jaafari et al. (2019) in the Zagros ecoregion of western Iran, they used three climate conditions: annual temperature, rainfall, and wind effects. However, the results showed that only the annual temperature parameter was relatively important. Similarly, Naderpour et al. (2020) studied the northern beaches of New South Wales, Australia, using three climate factors: annual rainfall, annual humidity, and annual wind speed. Holdrege et al. (2024) studied the Great Basin sagebrush regions across 13 western U.S. states and used three climate factors: annual mean temperature, annual precipitation, and the proportion of summer (June to August) precipitation to total annual precipitation. The study found that areas with low summer precipitation proportion, medium to high annual precipitation, and high temperatures had the highest wildfire probability. Similarly, Aldersley et al. (2011) used annual cumulative precipitation, annual lightning frequency, monthly precipitation frequency, monthly average temperature, and climate data from the British Atmospheric Data Centre (BADC) to study the drivers of monthly fire area on a global and regional scale. They found that climate factors such as high temperatures, moderate precipitation, and drought were the most important drivers globally. Mansuy et al. (2019) conducted a study across the U.S. and Canada. They found that climate variables were the primary control factors controlling burned area in most ecoregions, both inside and outside protected areas, surpassing landscape and anthropogenic factors. The climate data used in their study included annual average precipitation, annual average relative humidity, the Hargreaves climate moisture deficit (an annual measure of energy and moisture), and degree days below 18°C. Abdollahi and Pradhan (2023) conducted a wildfire prediction study in the Gippsland region of Victoria, Australia, and found that humidity, wind speed, and rainfall were the most influential factors in wildfire prediction. Zhao and Liu (2021) used data from over 600 meteorological stations in forest, grassland, and shrubland areas to calculate two drought indices, the Keetch-Byram Drought Index and the Standardized Precipitation Index. They then used global-scale PDSI data to mask these indices and assess long-term wildfire risk in China. The results showed that drought indices had a stronger correlation with seasonal and annual fire occurrences compared to climate indicators such as temperature and precipitation.